Circular No. 6406

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996ac IN ANONYMOUS GALAXY R. H. McNaught reports his discovery of an apparent supernova on

a U film taken on May 21.4 UT by K. S. Russell with the U.K. Schmidt Telescope, at which time SN 1996ac was fainter by ∼ 1 mag than on May 9.5, when its magnitude was ∼ 16.5 on a similar exposure taken by M. J. Drinkwater. SN 1996ac is located at α = 12^{h}24^{m}11^{s}.24, δ = +1^{o}24′14′′.7 (equinox 2000.0), which is 2′′.2 west and 3′′.2 north of the galaxy’s center. A nearby star (mag ∼ 17.5) has position end figures 13^{s}.78, 23′46′′.6. No star appears at the position of SN 1996ac on an I survey plate taken on 1991 Apr. 6, nor is anything present on a Palomar Sky Survey B film taken on 1992 Apr. 7.

L. Wang and J. C. Wheeler, University of Texas, report: “A lowresolution (FWHM 1.2 nm) spectrogram of SN 1996ac was obtained on May 23 UT using the 2.1-m telescope at McDonald Observatory. The spectrum confirms that this is a type-Ia supernova about a month past optical maximum.”

GRO J1655–40 K. Horne, E. T. Harlaftis, and R. Baptista, University of St. Andrews;

C. Hellier and A. Allan, Keele University; H. Johnston, Anglo-Australian Observatory; J. Patterson and J. Kemp, Columbia University; C. Haswell, University of Sussex and Columbia University; and W. Chen, Goddard Space Flight Center, report on the optical/ultraviolet brightening of GRO J1655–40 during the recent x-ray outburst (IAUC 6393): “Cerro Tololo photometry on May 10.41 UT gave Johnson magnitudes U = 17.0, B = 16.8, V = 15.4, R = 14.4, I = 13.4, which fit a power-law spectrum with slope 1/3, reddened by E(B −V) = 1.5 mag. Anglo-Australian Telescope spectra obtained on May 10.48, 11.57, and 12.57 confirm the red continuum slope, with emission from He II at 468.6 nm and CIII/NIII at 464.0 nm (FWHM = 2 nm, EW = 0.2 nm), and broad absorption filled in by multi-component emission at Hβ and Hα. Simultaneous HST/XTE data on May 13.9–14.3 show 5- to 10-percent variations on timescales of 2 s–10 min. Deep 220-nm absorption in the HST spectrum suggests E(B −V) = 1.3 mag. HST/XTE observations were planned for May 20.59–20.80, 27.8–28.1, June 8.53, and two additional times in June.”

1996 May 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green